OpenStudy (anonymous):

Subtract the following Fractions. -7/8-1/6

OpenStudy (anonymous):

Hi, do you know what the first step to solving this problem might be?

OpenStudy (anonymous):

find a common denominator

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Yes, do you know what that might be?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

So, we'll have a common denominator of 24 if we multiply 8 by 3, and 6 by 4. Understand so far?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright, so because we multiplied the denominators, we must do the same to the numerators. 7*3=? 1*4=?

OpenStudy (anonymous):

21 and 4

OpenStudy (anonymous):

17/24

OpenStudy (anonymous):

Yes, so if we make our new equivalent fractions we'll have.\[-\frac{ 21 }{ 24 }-\frac{ 4 }{ 24 }\]Subtract the numerators:\[-21-4=-25\]\[-\frac{ 25 }{ 24 }=1\frac{ 1 }{ 24 }\]Got it, or more explanation?

OpenStudy (anonymous):

Put a negative in front of that mixed fraction. My bad.

OpenStudy (anonymous):

\[-1\frac{ 1 }{ 24 }\]

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Blaize had 5/8 of a pound of turkey left over after a meal. He used 1/4 pount of it to make sandwiches for his family the next day. How many pounds of turkey are left? *

OpenStudy (anonymous):

So, common denominator. 4*2=8 1*2=2 New fraction:\[\frac{ 2 }{ 8 }\]Subtract:\[\frac{ 5 }{ 8 }-\frac{ 2 }{ 8 }\]\[5-2=3\]\[\frac{ 3 }{ }\]

OpenStudy (anonymous):

\[\frac{ 3 }{ 8 }\]The '8' didn't show up for whatever reason.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

what grade are u in?

OpenStudy (anonymous):

11th

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

im in 7th

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok yay

OpenStudy (anonymous):

Add the following fractions -2 3/8+(-4 1/2)

OpenStudy (anonymous):

In the second mixed fraction, we need to make the '2' an '8', or our common denominator. 2*4=8 4/4=4 New fraction:\[\frac{ 4 }{ 8 }\]New equation:\[-2\frac{ 3 }{ 8 }+(-4\frac{ 4 }{ 8 })\]\[-2+(-4)=-6\]\[-3+(-4)=-7\]\[-6\frac{ 7 }{ 8 }\]